A study is to be conducted to determine customer preference for short-travel push buttons on an automobile radio. There are two types: tact and rubber. The tact push buttons have a hard, firm feel, whereas the rubber push buttons have a soft, smooth feel. 10 people will be random selected and divided into two groups. Each group will be asked to rate one type of button on a scale from 1 to 10, with 10 denoting the highest level of satisfaction with performance of the push buttons.

A. Prepare a data layout.

B. Write a mathematical model for this experiment.

C. Prepare a partial ANOVA summary. Include the expected mean squares. Determine the appropriate tests that can be run and comment.
Answer #1

Solution

Let xij represent the jth observation in the ith row, j = 1,2,…,n; i = 1,2,……,r

Then the ANOVA model is: xij = µ + αi + εij, where µ = common effect, αi = effect of ith row, and εij is the error component which is assumed to be Normally Distributed with mean 0 and variance σ2.

Null hypothesis: H0: α1 = α2 = α3 = 0 Vs Alternative: HA: H0 is false [at least one αi is different from others]

Part (A)

Button Type

Ratings

Tact

x11, x12,…….., x1n1

Rubber

X21, x22,…….., x2n2

xij represents the rating of the jth person for the ith type of button, j = 1, 2, ….., ni, i = 1, 2; n1 = number of persons in the group rating Tact and n2 = number of persons in the group rating Rubber. ANSWER

Part (b)

The mathematical model is: xij = µ + αi + εij, where µ = common effect, αi = effect of ith type of button, and εij is the error component which is assumed to be Normally Distributed with mean 0 and variance σ2.

Null hypothesis: H0: α1 = α2 Vs Alternative: HA: H0 is false [α1 is different from α2]

ANSWER

Part (c)

ANOVA TABLE      

Source of

Variation

Degrees of Freedom (DF)

Sum of squares (SS)

Mean Sum

of squares

(MS = SS/DF)

Fobs

Fcrit*

Significance**

Row

r - 1

SSR

MSR/MSE

Error

rn - r

SSE

Total

rn - 1

SST

NOTE:

     * Fcrit: upper α% point of F-Distribution with degrees of freedom n1 and n2, where n1

     is the DF for Row and n2 is the DF for Error

     ** Significance: Fobs is significant if Fobs > Fcrit

Now, to work out the solution,

Terminology:

Row total = xi.= sum over j of xij

Grand total = G = sum over i of xi.

Correction Factor = C = G2/N, where N = total number of observations = r x n =

Total Sum of Squares: SST = (sum over i,j of xij2) – C

Row Sum of Squares: SSR = {(sum over i of xi.2)/(n)} – C

Error Sum of Squares: SSE = SST – SSR

Mean Sum of Squares = Sum of squares/Degrees of Freedom

DONE